OpenStudy (anonymous):

Solve: –6 + 3x = –9 Select one: a. –1 b. –6 c. –5 d. –3

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Are we solving for x?

OpenStudy (anonymous):

idk it just says solve

OpenStudy (anonymous):

i think its A

OpenStudy (anonymous):

Yeah it only makes sense if we are, alright so first step is to add both sides by 6 than divide by 3. \[3x=-9+6 \implies 3x = -3\] \[x = \frac{ -3 }{ 3} \implies-1\]

OpenStudy (anonymous):

This is a simple order of operations equation.

OpenStudy (anonymous):

Solve: x + 9 = 5(4x – 2) Select one: A. 11/19 B.–1 c. 1 d.-1/19

OpenStudy (anonymous):